# Simultaneous Pick and Place Detection by Combining SE(3) Diffusion Models with Differential Kinematics

### v1

@@ III-B Denoising Process Constrained by Differential IK
This section introduces a denoising process constrained by the robot kinematics (see Fig. 2 bottom for illustration). For compact notation, we denote the cases of pick and place by subscription $a$ and $b$, respectively, and $i\in\{a,b\}$ to represent both cases. For simplicity, we first discuss the pick-only case. We start from a heuristic collision-free initial joint position $\bm{q}_{a}(k=K)\in\mathbb{R}^{n}$ where $n$ denotes the number of joints of the robot, and iteratively update the joint position by $\bm{q}_{a}(k-1)=\bm{q}_{a}(k)+\dot{\bm{q}}_{a}(k)\Delta t$. Note that here the timestep $k$ is for conditioning the diffusion model (in this work we set $K=100$) and thus does not have physical meaning, while time delta $\Delta t$ has a unit of second to convert velocity to displacement (in this work, we set $\Delta t=1$). Both $K$ and $\Delta t$ need to be consistent with the one used in III-A. In the following discussions we omit $k$ for simple notation. At each timestep, we solve the forward kinematics to acquire the hand pose ${}_{h}T_{a}\in\mathbb{R}^{4\times 4}$ ($h$ stands for “hand”). This is fed into the diffusion model together with the timestep $t$, which outputs the target hand velocity ${}_{h}\bm{\xi}_{a}$. This leads to a simple quadratic programming (QP) based Diff-IK [31]:
$\min_{\dot{\bm{q}}_{a}}||_{h}\bm{\xi}_{a}-J_{a}\dot{\bm{q}}_{a}||^{2}~{}~{}s.t.$ (1)
$\bm{q}_{\mathrm{min}}-\bm{q}_{a}\leq\dot{\bm{q}}_{a}\Delta t\leq\bm{q}_{\mathrm{max}}-\bm{q}_{a}$
${}_{c}J_{a}\dot{\bm{q}}_{a}\leq 0.$
@@
The first constraint denotes the joint limit, with $\bm{q}_{\mathrm{min}},\bm{q}_{\mathrm{max}}$ being the lower and upper limits. We apply the second constraint only when any of the robot links are closer to the environment than a threshold, where ${}_{c}J\in\mathbb{R}^{m\times n}$ is the contact jacobian with $m$ denoting the number of contacts.
@@
To consider the placement, we introduce the placement joint position $\bm{q}_{b}\in\mathbb{R}^{n}$ and iteratively update it by computing $\dot{\bm{q}}_{b}$. At each time step, we get the placed object pose ${}_{o}T_{b}\in\mathbb{R}^{4\times 4}$ ($o$ stands for “object”) so that ${}_{h}T_{b}^{-1}{{}_{o}T_{b}}={{}_{h}T_{a}^{-1}}{{}_{o}T_{a}}$ stands. Here, ${}_{o}T_{a}$ is the pose of the target object from the observation. This formulates that the relative transformation between the hand and the object must be consistent for pick and place because we assume that the object doesn’t move in the hand. In order to let ${}_{o}T_{b}$ converge to the target place pose ${}_{o}T_{b}{{}^{\mathrm{target}}}$, we compute the target place object velocity ${}_{o}\bm{\xi}_{b}=[_{o}\bm{w}_{b}^{\top},{{}_{o}\bm{v}}_{b}^{\top}]^{\top}$ based on their pose difference. If multiple choices of place pose were given, ${}_{o}\bm{\xi}_{b}$ can be selected as the minimum norm one. By adding another decision variable $\bm{\xi}_{b}=[\bm{w}_{b}^{\top},{\bm{v}}_{b}^{\top}]^{\top}$, the problem in Eq. 1 is now extended as
$\min_{\dot{\bm{q}}_{i},\bm{\xi}_{b}}||_{h}\bm{\xi}_{a}-J_{a}\dot{\bm{q}}_{a}||^{2}+({{}_{o}\bm{\xi}}_{b}-\bm{\xi}_{b})^{\top}Q_{b}({{}_{o}\bm{\xi}}_{b}-\bm{\xi}_{b})~{}~{}s.t.$ (2)
${{}_{o}R}_{a}^{\top}J_{a}^{w}\dot{\bm{q}}_{a}={{}_{o}R}_{b}^{\top}(J_{b}^{w}\dot{\bm{q}}_{b}-\bm{w}_{b})$
${{}_{o}R}_{a}^{\top}J_{a}^{v}\dot{\bm{q}}_{a}={{}_{o}R}_{b}^{\top}(J_{b}^{v}\dot{\bm{q}}_{b}-\bm{v}_{b}-\bm{w}_{b}\times({{}_{h}\bm{p}}_{b}-{{}_{o}\bm{p}}_{b}))$
@@
Here, $Q_{b}\in\mathrm{diag}(6)$ is placing weight matrix. If placing with any rotation around the z-axis is accepted, e.g., uprighting a bottle, we can set $Q_{b}[3,3]=0$. Similarly, we can also allow arbitrary horizontal translation inside a region, e.g., anywhere on a table, by setting $Q_{b}[4,4]=Q_{b}[5,5]=0$, with an additional inequality constraint on $\bm{\xi}_{b}$ similar to the joint limit constraint. The placing joint velocity $\dot{\bm{q}}_{b}$ only appears in the constraints. The first and second constraints require the relative velocity between the hand and the object to be identical for pick and place in order to maintain that the relative hand-object transform for pick and place are the same at all timesteps. The equation is asymmetric because ${{}_{o}T}_{a}$ is constant. Note that $R\in\mathbb{R}^{3\times 3},\bm{p}\in\mathbb{R}^{3}$ is the rotational and translational part of $T$, and $J^{w}_{i},J^{v}_{i}\in\mathbb{R}^{3\times n}$ is the upper and lower half of $J_{i}$.
@@
While there are a large number of works in motion planning, they typically assume all obstacles are known. However, this is not the case when grasping unknown objects placed in a cluttered way based on partial observation. We therefore take a heuristic approach where the hand linearly moves down from ${{}^{e}_{h}T}$ to ${{}^{a}_{h}T}$ and moves forward from ${{}^{a}_{h}T}$ to ${{}_{h}T}$. After closing the hand, we lift it vertically to the same height as ${{}^{e}_{h}T}$, denoting it as ${{}^{l}_{h}T}$. A dense joint trajectory is then acquired by linearly interpolating ${{}^{e}_{h}T},{{}^{a}_{h}T},{{}_{h}T},{{}^{l}_{h}T}$ in the Cartesian space and solving Diff-IK. We also apply this process for the placement, in reverse order. As ${{}^{l}_{h}T}$ is relatively safe from unobserved obstacles, we use a PRM [32]-based motion planner to connect between ${{}^{l}_{h}T_{a}}$ and ${{}^{l}_{h}T_{b}}$.
The overall denoising process is summarized as follows. We start from multiple random $\bm{q}_{i}(K)$ based on heuristics. We use $\bm{q}_{i}(K)$ as the initial value of ${}^{a}\bm{q}_{i}(K),{{}^{e}\bm{q}_{i}(K)}$ also. At each timestep $k$, we solve the forward kinematics to get the hand pose ${{}_{h}T_{a}}(k)$, which is fed into the network. Based on the network output ${{}_{h}\bm{\xi}_{a}(k)}$, we solve the optimization problem to get $\dot{\bm{q}}_{i}(k),{{}^{a}\dot{\bm{q}}_{i}(k)},{{}^{e}\dot{\bm{q}}_{i}(k)}$ to iteratively update the joint configuration. After the iterations from $k=K$ to $k=0$, we project ${{}_{h}T_{a}}(0)$ to the grasp validity volume output by the network to filter out invalid grasps, and use the gravity rejection score volume to sort for the best one. Finally, we compute the dense joint trajectory and send it to the robot.
@@ IV Experiment and Evaluation
This section evaluates our approach using simulation and real hardware with the same problem setting. A robotiq 2F-85 gripper [35] is attached to a Franka-Emika Panda Robot [36] located in front of a table. In each scene, one to five unknown household objects are randomly arranged inside a 30 cm square region on the table surface. In each trial, the robot picks one object and places it upright on a small rack, i.e., we allow arbitrary rotation around the z-axis. We follow the metrics proposed in [9] where the success ratio (SR) represents the number of successful grasps divided by the number of total trials, and the clear ratio (CR) represents the number of successful grasps divided by the total number of objects. However, since we also target placement, we count a trial as success only if the object is placed upright with less than 5 cm position error. Similarly to the case of [9], each scene is terminated after two consecutive pick-place or detection failures.
@@ IV-A Comparison with Baseline
To our best knowledge, there are no prior works that share the same problem setting with us. We therefore create a baseline based on a naive depth-first search pipeline. We select [12] as the grasp detector because our scene encoder and decoder are almost the same as the one in [12], and it also considers power grasping by the Robotiq gripper. This feedforward grasp detector detects all feasible grasp poses in the scene in the gravity rejection score order. For each of the grasp poses, we first check whether it grasps the target object by checking the intersection between the hand region with the masked point cloud. We then try to create the pick trajectory based on the interpolation of ${{}^{e}_{h}T_{a}},{{}^{a}_{h}T_{a}},{{}_{h}T_{a}},{{}^{l}_{h}T_{a}}$, where ${{}_{h}T_{a}}$ is the detection output. Since we allow arbitrary yaw rotation for placing, we divide the yaw by 12. For each of the placement candidates, we try to interpolate place trajectory by ${{}^{l}_{h}T_{b}},{{}_{h}T_{b}},{{}^{a}_{h}T_{b}},{{}^{e}_{h}T_{b}}$, where ${{}_{h}T_{b}}$ is derived so that the relative transformation between the hand and the object is kept for pick and place. If this process fails at any stage, we move to the next grasp pose until we find the first feasible trajectory for the robot to execute.
@@
Diff. F.F. Diff. F.F. Diff. F.F.
@@ IV-C Real Robot Verification
  Baseline Ours Ours w/ pre-place refinement
@@
Table III summarizes the result, which is consistent with the simulation experiment where our approach outperformed the baseline in both SR/CR and computation time. The computation times are longer because they include the object detection and pose estimation pipeline, while the 6-second reduction from the baseline to our approach is due to our contribution. During the experiment, we observed that pose estimation error was one of the major sources of placement failure. We therefore introduced a variant of our approach: we used another pair of cameras located near the placement location to re-identify the object’s pose before the robot placed it. We used the updated ${}_{o}T_{b}$ to solve the place trajectory again, which turned out to effectively compensate for the pose estimation error and the error due to the object’s movement inside the hand. We show its result in the rightmost column of the table. Notice that the time for the refinement is not included in the table.
@@ V Limitation and Future Work
During the evaluations, we observed two major failure modes. The first one is grasp failure due to inaccurate approach direction. Since both the grasp validness volume and the gravity rejection score volume are based on the voxel location, they can only verify the grasp’s translation but not its rotation. Introducing an additional post-processor to evaluate SE(3) grasp poses will improve performance while it remains in our future works. The other failure case is placement failure: if the object moves inside the hand during the pick-to-place trajectory, there is a high chance that it will fall over during the placement. While the gravity rejection score represents the magnitude of gravity direction disturbance a grasp can resist, the gravity direction relative to the grasp varies during the pick-to-place trajectory execution. Extending it to consider multiple disturbance directions also remains in our future works.

### v2

@@ III-B Denoising Process Constrained by Differential IK
This section introduces a denoising process constrained by the robot kinematics (see Fig. 2 bottom for illustration). For compact notation, we denote the cases of pick and place by subscription $\alpha$ and $\beta$, respectively, and $i\in\{\alpha,\beta\}$ to represent both cases. For simplicity, we first discuss the pick-only case. We start from a heuristic collision-free initial joint position $\bm{q}_{\alpha}(k=K)\in\mathbb{R}^{n}$ where $n$ denotes the number of joints of the robot, and iteratively update the joint position by $\bm{q}_{\alpha}(k-1)=\bm{q}_{\alpha}(k)+\dot{\bm{q}}_{\alpha}(k)\Delta t$. Note that here the timestep $k$ is for conditioning the diffusion model (in this work we set $K=100$) and thus does not have physical meaning, while time delta $\Delta t$ has a unit of second to convert velocity to displacement (in this work, we set $\Delta t=1$). Both $K$ and $\Delta t$ need to be consistent with the one used in III-A. In the following discussions we omit $k$ for simple notation. At each timestep, we solve the forward kinematics to acquire the hand pose $_{h}T_{\alpha}\in\mathbb{R}^{4\times 4}$ ($h$ stands for “hand”). This is fed into the diffusion model together with the timestep $t$, which outputs the target hand velocity $_{h}\bm{\xi}_{\alpha}$. This leads to a simple quadratic programming (QP) based Diff-IK [31]:
$\min_{\dot{\bm{q}}_{\alpha}}||_{h}\bm{\xi}_{\alpha}-J_{\alpha}\dot{\bm{q}}_{\alpha}||^{2}~~s.t.$ (1)
$\bm{q}_{\mathrm{min}}-\bm{q}_{\alpha}\leq\dot{\bm{q}}_{\alpha}\Delta t\leq\bm{q}_{\mathrm{max}}-\bm{q}_{\alpha}$
$_{c}J_{\alpha}\dot{\bm{q}}_{\alpha}\leq 0.$
@@
The first constraint denotes the joint limit, with $\bm{q}_{\mathrm{min}},\bm{q}_{\mathrm{max}}$ being the lower and upper limits. We apply the second constraint only when any of the robot links are closer to the environment than a threshold, where $_{c}J\in\mathbb{R}^{m\times n}$ is the contact jacobian with $m$ denoting the number of contacts. Note that the “robot” includes the hand also, and the “environment” includes primitive shapes extraced from the initial observation (in this work the table top) and point cloud of the objects aquired at each observation.
@@
To consider the placement, we introduce the placement joint position $\bm{q}_{\beta}\in\mathbb{R}^{n}$ and iteratively update it by computing $\dot{\bm{q}}_{\beta}$. At each time step, we get the placed object pose $_{o}T_{\beta}\in\mathbb{R}^{4\times 4}$ ($o$ stands for “object”) so that $_{h}T_{\beta}^{-1}{{}_{o}T_{\beta}}={{}_{h}T_{\alpha}^{-1}}{{}_{o}T_{\alpha}}$ stands. Here, $_{o}T_{\alpha}$ is the pose of the target object from the observation. This formulates that the relative transformation between the hand and the object must be consistent for pick and place because we assume that the object doesn’t move in the hand. In order to let $_{o}T_{\beta}$ converge to the target place pose $_{o}T_{\beta}{{}^{\mathrm{target}}}$, we compute the target place object velocity $_{o}\bm{\xi}_{\beta}=[_{o}\bm{w}_{\beta}^{\top},{{}_{o}\bm{v}}_{\beta}^{\top}]^{\top}$ based on their pose difference. If multiple choices of place pose were given, $_{o}\bm{\xi}_{\beta}$ can be selected as the minimum norm one. By adding another decision variable $\bm{\xi}_{\beta}=[\bm{w}_{\beta}^{\top},{\bm{v}}_{\beta}^{\top}]^{\top}$, the problem in Eq. 1 is now extended as
$\min_{\dot{\bm{q}}_{i},\bm{\xi}_{\beta}}||_{h}\bm{\xi}_{\alpha}-J_{\alpha}\dot{\bm{q}}_{\alpha}||^{2}+({{}_{o}\bm{\xi}}_{\beta}-\bm{\xi}_{\beta})^{\top}Q_{\beta}({{}_{o}\bm{\xi}}_{\beta}-\bm{\xi}_{\beta})~~s.t.$ (2)
${{}_{o}R}_{\alpha}^{\top}J_{\alpha}^{w}\dot{\bm{q}}_{\alpha}={{}_{o}R}_{\beta}^{\top}(J_{\beta}^{w}\dot{\bm{q}}_{\beta}-\bm{w}_{\beta})$
${{}_{o}R}_{\alpha}^{\top}J_{\alpha}^{v}\dot{\bm{q}}_{\alpha}={{}_{o}R}_{\beta}^{\top}(J_{\beta}^{v}\dot{\bm{q}}_{\beta}-\bm{v}_{\beta}-\bm{w}_{\beta}\times({{}_{h}\bm{p}}_{\beta}-{{}_{o}\bm{p}}_{\beta}))$
@@
Here, $Q_{\beta}\in\mathrm{diag}(6)$ is placing weight matrix. If placing with any rotation around the z-axis is accepted, e.g., uprighting a bottle, we can set $Q_{\beta}[3,3]=0$. This problem is typically termed an incomplete orientation constraint [32]. Similarly, we can also allow arbitrary horizontal translation inside a region, e.g., anywhere on a table, by setting $Q_{\beta}[4,4]=Q_{\beta}[5,5]=0$, with an additional inequality constraint on $\bm{\xi}_{\beta}$ similar to the joint limit constraint. The placing joint velocity $\dot{\bm{q}}_{\beta}$ only appears in the constraints. The first and second constraints require the relative velocity between the hand and the object to be identical for pick and place in order to maintain that the relative hand-object transform for pick and place are the same at all timesteps. The equation is asymmetric because ${{}_{o}T}_{\alpha}$ is constant. Note that $R\in\mathbb{R}^{3\times 3},\bm{p}\in\mathbb{R}^{3}$ is the rotational and translational part of $T$, and $J^{w}_{i},J^{v}_{i}\in\mathbb{R}^{3\times n}$ is the upper and lower half of $J_{i}$.
@@
While there are a large number of works in motion planning, they typically assume that all obstacles are known. However, this is not the case when grasping unknown objects placed in a cluttered way based on partial observation. We therefore take a more conservative approach where the hand linearly moves down from $^{e}_{h}T$} to $^{a}_{h}T$} and moves forward from $^{a}_{h}T$} to $_{h}T$}. After closing the hand, we lift it vertically to the same height as $^{e}_{h}T$}, denoting it as $^{l}_{h}T$}. A dense joint trajectory is then acquired by linearly interpolating ${{}^{e}_{h}T},{{}^{a}_{h}T},{{}_{h}T},{{}^{l}_{h}T}$ in the Cartesian space and solving Diff-IK. We also apply this process for the placement, in reverse order. As $^{l}_{h}T$} is relatively safe from unobserved obstacles, we use a PRM [33]-based motion planner to connect between $^{l}_{h}T_{\alpha}$} and $^{l}_{h}T_{\beta}$}.
The overall denoising process is summarized as follows. We start from multiple random $\bm{q}_{i}(K)$ based on heuristics. We use $\bm{q}_{i}(K)$ as the initial value of $^{a}\bm{q}_{i}(K),{{}^{e}\bm{q}_{i}(K)}$ also. At each timestep $k$, we solve the forward kinematics to get the hand pose ${{}_{h}T_{\alpha}}(k)$, which is fed into the network. Based on the network output $_{h}\bm{\xi}_{\alpha}(k)$}, we solve the optimization problem to get $\dot{\bm{q}}_{i}(k),{{}^{a}\dot{\bm{q}}_{i}(k)},{{}^{e}\dot{\bm{q}}_{i}(k)}$ to iteratively update the joint configuration. After the iterations from $k=K$ to $k=0$, we project ${{}_{h}T_{\alpha}}(0)$ to the grasp validity volume output by the network to filter out invalid grasps, and use the gravity rejection score volume to sort for the best one. Finally, we compute the dense joint trajectory and send it to the robot.
@@ IV Experiment and Evaluation
This section evaluates our approach using simulation and real hardware with the same problem setting. A robotiq 2F-85 gripper [36] is attached to a Franka-Emika Panda Robot [37] located in front of a table. In each scene, one to five unknown household objects are randomly arranged inside a 30 cm square region on the table surface. In each trial, the robot picks one object and places it upright on a small rack, i.e., we allow arbitrary rotation around the z-axis. We follow the metrics proposed in [9] where the success ratio (SR) represents the number of successful grasps divided by the number of total trials, and the clear ratio (CR) represents the number of successful grasps divided by the total number of objects. However, since we also target placement, we count a trial as success only if the object is placed upright with less than 5 cm position error.111 When the placement is successful, the position error is tyically less than 1 or 2 cm. The tolerance of 5 cm here is simply to reject cases where the object is dropped before placing. Similarly to the case of [9], each scene is terminated after two consecutive pick-place or detection failures.
@@ IV-A Comparison with Baseline
To our best knowledge, there are no prior works that share the same problem setting with us. We therefore create a baseline based on a naive depth-first search pipeline. We select [12] as the grasp detector because our scene encoder and decoder are almost the same as the one in [12], and it also considers power grasping by the Robotiq gripper. This feedforward grasp detector detects all feasible grasp poses in the scene in the gravity rejection score order. For each of the grasp poses, we first check whether it grasps the target object by checking the intersection between the hand region with the masked point cloud. We then try to create the pick trajectory based on the interpolation of ${{}^{e}_{h}T_{\alpha}},{{}^{a}_{h}T_{\alpha}},{{}_{h}T_{\alpha}},{{}^{l}_{h}T_{\alpha}}$, where $_{h}T_{\alpha}$} is the detection output. Since we allow arbitrary yaw rotation for placing, we divide the yaw by 12. For each of the placement candidates, we try to interpolate place trajectory by ${{}^{l}_{h}T_{\beta}},{{}_{h}T_{\beta}},{{}^{a}_{h}T_{\beta}},{{}^{e}_{h}T_{\beta}}$, where $_{h}T_{\beta}$} is derived so that the relative transformation between the hand and the object is kept for pick and place. If this process fails at any stage, we move to the next grasp pose until we find the first feasible trajectory for the robot to execute.
@@
Diff. $F.$ Diff. $F.$ Diff. $F.$
@@ IV-C Real Robot Verification
Table III summarizes the result, which is consistent with the simulation experiment where our approach outperformed the baseline in both SR/CR and computation time. The computation times are longer because they include the object detection and pose estimation pipeline, while the 6-second reduction from the baseline to our approach is due to our contribution. During the experiment, we observed that pose estimation error was one of the major sources of placement failure. We therefore introduced a variant of our approach: we used another pair of cameras located near the placement location to re-identify the object’s pose before the robot placed it. We used the updated $_{o}T_{\beta}$ to solve the place trajectory again, which turned out to effectively compensate for the pose estimation error and the error due to the object’s movement inside the hand. We show its result in the rightmost column of the table. Notice that the time for the refinement is not included in the table.
@@
  Baseline Ours
Ours w/
pre-place refinement
@@ V Limitation and Future Work
V-1 Failure Modes
During the evaluations, we observed two major failure modes. The first one is grasp failure due to inaccurate approach direction. Since both the grasp validness volume and the gravity rejection score volume are based on the voxel location, they can only verify the grasp’s translation but not its rotation. Introducing an additional post-processor to evaluate SE(3) grasp poses will improve performance while it remains in our future works. The other failure case is placement failure: if the object moves inside the hand during the pick-to-place trajectory, there is a high chance that it will fall over during the placement. While the gravity rejection score represents the magnitude of gravity direction disturbance a grasp can resist, the gravity direction relative to the grasp varies during the pick-to-place trajectory execution. Extending it to consider multiple disturbance directions is a possible extension.
V-2 Local Minima at Denoising Process
As another limitation, we occasionally observed that the hand was trapped in the concave-shaped part of the surrounding obstacles during denoising, especially when the scene is heavily cluttered. This is because our collision avoidance is based on hard inequality constraints. Relaxing the constraints by introducing potential fields or first reaching the hand without considering the collision and then repelling the colliding part from the obstacles may imporve the detection success ratio, while it still remains as future work.
V-3 Computation Efficiency
In this work, we implemented the kinematics loop in Python for the ease of combining with the diffusion model. In order to align with this, we also implemented the baseline naively in Python, which is less performant than the state-of-the-art libraries such as MoveIt! [44]. While we believe this is an orthogonal to our contribution, i.e., we can also introduce such optimization technics to our approach, actual validation still remains as our future work.